If the npv of a shop is R195000 and the profitability index is 1,24375, the initial investment in shop is?
1
Expert's answer
2021-07-28T16:43:14-0400

"PI=(NPV+" initial investment)"\div" initial investment

"1.24375=(195 000+Initial \space investment)\div""initial\space investment"

"1.24375 \times initial \space investment=195000+initial\space investment"

"1.24375\times initial\space investment-initial\space investment=195000"

"initial\space investment (1.24375 -1)=195000,\\ initial \space investment=195000/0.24374=800032.82"

the initial investment in shop is R800032.82